OpenStudy (anonymous):

solve the equation using square roots. x2-9=0

OpenStudy (anonymous):

(x-3)(x+3)=0 x=3, -3

OpenStudy (anonymous):

(x-3)(x+3)

OpenStudy (anonymous):

factors to (x+3)(x-3)=0

OpenStudy (anonymous):

find all values of x that make the answer 0